# MMSearch-Plus: Benchmarking Provenance-Aware Search for Multimodal Browsing Agents

### v2

@@ Abstract
Project page: https://mmsearch-plus.github.io/
@@ 1 Introduction
Recent advances in large multimodal language models (MLLMs) have enabled them to act as capable agents that integrate language, vision, and tool use to accomplish complex tasks. A growing body of work studies how MLLMs leverage external search engines—both text- and image-based—to answer information-seeking questions.
@@
Multimodal browsing benchmarks such as MMSearch (Jiang et al., 2024) primarily test a model’s ability to coordinate text and images. However, many MMSearch-style questions can be solved by relatively fixed workflows that require little genuine multimodal reasoning. In particular, MMSearch heavily relies on external image search: the MLLM often plays the role of an orchestrator rather than a visual reasoner. When the search engine retrieves highly relevant images, even a unimodal LLM—without visual perception—can frequently answer by reasoning over the accompanying page text alone. This occurs because a single strong image search can surface pages whose surrounding text already contains the answer. In such cases, the image search tool and the MLLM are partially interchangeable as sources of image-related information; the only distinctly multimodal step is often image–source cross-validation, a narrow slice of multimodal reasoning.
Empirically, MMSearch requires short reasoning chains and achieves relatively high model scores: GPT-4o exceeds 60% on MMSearch (end-to-end VQA), and MMSearch-R1-7B reports 54.6% average accuracy (Wu et al., 2025). By contrast, recent text-only browsing benchmarks emphasize persistence and creative, multi-step search. BrowseComp (Wei et al., 2025) targets web navigation for hard-to-find, entangled information; on this task, GPT-4o and GPT-4.5 each score below 1% in direct-answer settings and under 2% even with browsing tools (Wei et al., 2025). “Deep Research” reports 51.5% on BrowseComp, though it was trained for the benchmark and public implementation details remain limited, while open-source agents such as WebSailor reach roughly 12% after substantial engineering (Li et al., 2025a). Taken together, these results suggest that current multimodal benchmarks may not yet match the difficulty of complex text-only browsing, motivating a BrowseComp-style multimodal benchmark with similarly demanding reasoning.
What makes a challenging benchmark for multimodal browsing agents.
Building on these insights, we argue that a strong multimodal benchmark should: (i) require fine-grained, exhaustive visual reasoning that compels models to mine subtle, localized cues rather than rely on a single salient (“dominant”) entity that occupies most of the image; and (ii) enforce source verification under retrieval noise—i.e., discriminating authentic sources when image results are noisy or conflicting and validating images embedded in webpages. Concretely, the benchmark should stress robust cross-modal evidence gathering, careful verification, and resilience to near-duplicates. For example, two football matches from adjacent seasons may share nearly identical visuals (kits, stadium angles). Correctly answering requires inspecting localized markers (e.g., patch details, sponsor variants, scoreboard typography) and cross-checking dates and venues against trustworthy sources. In real applications, difficult multimodal information-seeking tasks exhibit three recurring challenges that our benchmark targets:
@@
Noisy or conflicting retrieval. Image search may return non-relevant or contradictory results; agents must identify authentic sources via fine-grained multimodal understanding and cross-source consistency checks.
@@
Exhaustive, part-based visual reasoning. When holistic matching fails, agents must reason over parts and subregions—often iteratively cropping and re-searching—to piece together fragmented clues that identify the target entity or event.
@@
Professional multimodal handling and long, tool-augmented chains. Agents must extract answers from heterogeneous media (images, documents, videos) across the web, sustain long reasoning/tool-use chains, and invoke visual tools programmatically (e.g., cropping, OCR) when helpful, akin to advanced agents such as o3.
Our work, MMSearch-Plus, aims to bridge the gap by combining the persistence and high-reasoning demands of BrowseComp with truly multimodal workflows. Unlike (Geng et al., 2025), where images often collapse to a single identifiable entity, images in our benchmark contain multiple subtle visual signals that must be connected throughout the entire reasoning process. To systematically curate hard multimodal search questions, we introduce Spatial-Temporal Extrapolation, a principled curation procedure that seeds questions whose answers require extrapolating from spatial (localized visual cues, layouts, micro-text, uniforms, signage) and temporal signals (broadcast overlays, seasonal context, historical rosters/schedules) to identify events, dates, or locations not explicitly present in the prompt or image. This forces models to extract fine-grained cues, propagate them through iterative (image/text) searches, and verify provenance amidst retrieval noise—directly addressing the challenges above.
MMSearch-Plus targets realistic scenarios where answers cannot be read off from the prompt or the image alone. For example, given a screenshot from an esports broadcast, the agent may need to infer the match date—absent from both the query and the image—by integrating subtle overlays with external evidence. Similar patterns arise in sports, geolocation, and lifestyle vlogs, domains we explicitly include to move beyond typical web imagery.
Our key contributions can be summarized as follows:
We propose MMSearch-Plus, a challenging benchmark that integrates fine-grained multimodal reasoning (with sparse, localized visual signals) and external tool use, while maintaining the difficulty profile of text-based browsing tasks like BrowseComp (Wei et al., 2025).
We introduce the curation procedure Spatial-Temporal Extrapolation ( 2.1) to construct difficult multimodal search questions that require extrapolating from localized spatial cues and temporal signals to out-of-image facts.
We develop a general, model-agnostic search agent framework that interleaves text/image search with a Set-of-Mark zoom-in pathway ( 2.5). SoM supplies human-verified marks and region-seeded retrieval (zoom_in/image_search), enabling sustained fine-grained reasoning; we ablate SoM to quantify how cropping and thinking with images affect current MLLMs.
A concurrent work, BrowseComp-VL (Geng et al., 2025), increases difficulty mainly by expanding the text search space in a BrowseComp-like manner, while the image often reduces to a single identifiable entity that is straightforwardly searchable and quickly disposable. Another concurrent work, MM-BrowseComp (Li et al., 2025b), is related. Our contributions differ in (a) data sources and (b) a general search framework that also enables thinking with images via Set-of-Mark.
@@ 2.5 Search Agent Framework
3 Experiment
Evaluated Models.
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini) (Team & DeepMind, 2025), o3 (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a), representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen) (Bai et al., 2025). Expert human annotators were instructed to work with a Chrome browser for at least 10 minutes for each task, and capped at 20 minutes.
@@ 3 Experiment
To ensure a fair comparison, we standardize the search-and-reasoning pipeline across all models. We consider three search modes:
@@
Full Rollout. The model runs our search-agent framework and may call search tools as needed. Image-search settings match those in Image Search. For text search, we retrieve the top-5 websites by default. The process is capped at 20 search rounds (combining text and image search) to balance efficiency and coverage. In each text-search round, the model generates 3–5 refined queries; in each image-search round, it issues exactly one query. For Qwen, pilot studies showed that more than 10 rounds yield no further gains, so we cap its search at 10 rounds.
@@ 4 Results
Under No Search, open-source Qwen scored 0.0%, confirming that MMSearch+ resisted parametric shortcuts. Closed-source models remained far from saturation (GPT-5 10.3, Gemini-2.5-Pro 10.6, o3 15.1). With Image Search (o3 +4.2, GPT-5 +6.1, Gemini +5.8; Qwen +13.5), models gained from coarse disambiguation, yet multi-hop composition remained the main blocker. With iterative image and text search, Full Rollout produced large gains: o3 36.0 (+20.9 over no-search) and Gemini 23.8 (+13.2), while Qwen regressed to 6.1 (-7.4 over image-search). This indicates that the external information retrieved from search tools is essential to answering the questions correctly. Notably, easy split accuracy did not improve from Image Search to Full Rollout. We attribute this to two factors:
@@
Policy overhead and distractors. Full rollout added branching (tool choice, query rewriting). For easy items solvable by a single image search, extra hops often injected distractors (off-topic pages, SEO noise), increasing drift from the correct evidence.
Overall, rollout helped where progressive narrowing was necessary (academic, tech, media queries), but on easy items extra freedom reduced precision.
Adding SoM (Full Rollout + SoM) gave consistent, targeted gains: o3 37.7 (+1.7), Gemini 27.7 (+3.9), Qwen +1.0 over Full Rollout. Set-of-Mark anchored reasoning on fine-grained visual cues (micro-text, logos, key entities) before expanding the search space.
@@ 5.1 Observations
In Figure 5, all models made more text-search calls than image-search calls. In addition, MLLMs sometimes zoomed in on the SoM-cropped image subregion only to see it clearly, without making a subsequent image search using the sub-image as the query.
@@ 6 Related Work
Classical vision–language benchmarks probe compositional and commonsense reasoning in images and text (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021), with recent suites extending to long videos and interleaved modalities (Wu et al., 2024; Li et al., 2024; Hu et al., 2025; Jeong et al., 2025). Concurrent work on web agents studies planning, browsing, and evidence gathering in open worlds (Krishna et al., 2024; Wei et al., 2025; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025). Multimodal web search benchmarks begin to couple vision with browsing, but often emphasize entity matching over fine-grained visual reasoning (Jiang et al., 2024; Wu et al., 2025). Closest to our setting, MM-BrowseComp targets image/video evidence encountered during search and reveals substantial headroom even for frontier models (Li et al., 2025b). A complementary line promotes “thinking with images” via improved grounding and RL-style training to elicit image-first reasoning (Wu & Xie, 2024; Zheng et al., 2024; Fan et al., 2025; Zheng et al., 2025b; Su et al., 2025a; Xia et al., 2025). We differ by targeting sparse, spatial–temporal visual cues that must be transformed into targeted queries and multi-step plans, providing a focused stress test for multimodal browsing agents.
@@ 7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) reached only 15.1% without search and 36.0% with rollout, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieved 0.0% without search and 6.9% with rollout. These results indicate that current systems struggle to (i) reliably read and localize micro-signals (badges, micro-text, uniforms, layout geometry), (ii) decide when and how to crop and reuse visual evidence, and (iii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.

### v3

@@ 1 Introduction
Large multimodal language models (MLLMs) increasingly act as agents that integrate vision, language, and web search to answer information-seeking questions. Several recent benchmarks study this capability by pairing images with web browsing or image search tools. However, despite their multimodal interface, existing multimodal browsing benchmarks rarely demand sustained, fine-grained visual reasoning or long-horizon, tool-augmented search.
This paper introduces MMSearch-Plus, a new benchmark that closes this gap. MMSearch-Plus is designed to require (i) localized, exhaustive visual reasoning over subtle cues, (ii) robust verification under noisy or conflicting retrieval, and (iii) multi-step tool use that interleaves text and image search with region-level visual analysis. To support this, we contribute:
A challenging multimodal browsing benchmark that matches the long-horizon difficulty of text-based browsing tasks (e.g., BrowseComp) while adding visual reasoning requirements that cannot be bypassed by a single strong image search.
Spatial–Temporal Extrapolation, a principled curation procedure that constructs hard questions by requiring models to extrapolate from localized spatial cues and temporal signals to out-of-image facts such as dates, events, or locations.
A general, model-agnostic agent framework that interleaves text/image search with a Set-of-Mark (SoM) zoom-in pathway, enabling region-seeded retrieval (zoom_in/image_search) and systematic analysis of how cropping and “thinking with images” affect current MLLMs.
@@
We next motivate why such a benchmark is needed. Multimodal browsing benchmarks such as MMSearch (Jiang et al., 2024) primarily test an MLLM’s ability to coordinate text and images, but many of their questions admit relatively fixed workflows. A strong image search engine often retrieves pages whose surrounding text directly contains the answer, allowing unimodal LLMs—without visual perception—to perform competitively. In these cases, multimodality collapses to a narrow form of image–source cross-validation, and fine-grained visual reasoning plays only a minor role.
In contrast, recent text-only browsing tasks such as BrowseComp (Wei et al., 2025) emphasize persistence, multi-step evidence gathering, and complex search strategies. State-of-the-art MLLMs score below 2% with browsing tools, indicating substantial headroom and revealing a discrepancy: multimodal browsing benchmarks are far easier than their text-only counterparts, even though real-world multimodal tasks often require deeper reasoning.
Design principles for a challenging multimodal browsing benchmark.
Difficult web-based visual tasks typically exhibit three recurring challenges:
@@
Noisy or conflicting retrieval. Image search may surface near-duplicates, irrelevant scenes, or contradictory candidates; solving the task requires cross-source consistency checks and multimodal verification.
@@
Exhaustive, part-based visual reasoning. When holistic matches fail, agents must reason over subregions—cropping, zooming, and re-searching—to accumulate fragmented spatial cues.
@@
Long, tool-augmented reasoning chains. Realistic tasks require sustained coordination of visual tools (cropping, OCR), textual browsing, and multimodal evidence integration.
Benchmarks such as MMSearch do not consistently elicit these behaviors: many images contain a single dominant entity that quickly anchors a search, and the subsequent steps are primarily text-based.
MMSearch-Plus.
To construct tasks that inherently require fine-grained visual reasoning, MMSearch-Plus introduces Spatial–Temporal Extrapolation: we curate images containing multiple subtle signals—micro-text, uniform variations, broadcast overlays, layout signatures—that jointly imply facts not explicitly present in the prompt. Models must extract these cues, propagate them through iterative searches, and verify provenance amid retrieval noise. This better reflects real-world scenarios where neither the prompt nor the image alone suffices; for instance, determining the date of an esports broadcast or identifying a specific match from nearly identical sports frames. Finally, we develop a general agent framework with Set-of-Mark (SoM) to study how region-level visual grounding influences browsing agents. SoM provides human-verified marks and region-seeded queries, enabling “thinking with images” through structured zoom-in reasoning. Ablations quantify how cropping and localized visual extraction affect current MLLMs.
Together, MMSearch-Plus provides a high-difficulty, multimodal benchmark and analysis toolkit for evaluating the next generation of multimodal browsing agents. It differs from concurrent efforts such as BrowseComp-VL (Geng et al., 2025) and MM-BrowseComp (Li et al., 2025b) in both data sources and its agent framework, which emphasizes sustained fine-grained visual reasoning.
@@ 2.2 Data Collection
Dataset updates.
Some problems that originally required retrieval later became solvable without search because newer closed-source MLLMs began to recognize previously unfamiliar entities or figures. This temporal drift is unavoidable as models continue to incorporate more recent data. We therefore commit to regularly refreshing the benchmark to suppress such internal-knowledge shortcuts.
@@ 2.5 Search Agent Framework
3 Experiment
Evaluated Models.
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini) (Team and DeepMind, 2025), o3 (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a), representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen) (Bai et al., 2025). Expert human annotators were instructed to work with a Chrome browser for at least 10 minutes for each task, and capped at 20 minutes.
@@ 3 Experiment
To ensure a fair comparison, we standardize the search-and-reasoning pipeline across all models. We consider five search modes:
@@
Text Search. For text search, we retrieve the top-5 websites by default. The process is capped at 20 search rounds. In each text-search round, the model generates 3–5 refined queries
Full Rollout. The model runs our search-agent framework and may call search tools as needed. Image-search and text-search settings match those described before. The process is capped at 20 search rounds (combining text and image search) to balance efficiency and coverage. In each image-search round, it issues exactly one query. For Qwen, pilot studies showed that more than 10 rounds yield no further gains, so we cap its search at 10 rounds.
@@ 4 Results
Baseline performance.
Under No Search on the original MMSearch+ benchmark, open-source Qwen scored 0.0%, confirming that MMSearch+ resisted parametric shortcuts. Closed-source models remained far from saturation (GPT-5 10.3, Gemini-2.5-Pro 10.6, o3 15.1). With Image Search (o3 +4.2, GPT-5 +6.1, Gemini +5.8; Qwen +13.5), models gained from coarse disambiguation, yet multi-hop composition remained the main blocker. With iterative image and text search, Full Rollout produced large gains: o3 36.0 (+20.9 over no-search) and Gemini 23.8 (+13.2), while Qwen regressed to 6.1 (-7.4 over image-search). This indicates that the external information retrieved from search tools is essential to answering the questions correctly.
Performance drop on easy split.
Notably, easy split accuracy did not improve from Image Search to Full Rollout. For o3, we conducted a targeted error analysis on this split: there are 23 “easy” questions for which the image-only setting is correct but the full-rollout setting is not, corresponding to a 7.4-point absolute accuracy loss on our 311-task benchmark. We randomly sampled 10 of these 23 cases and manually inspected the trajectories. In 9 out of 10, the full-rollout trajectory never invoked image search at all: the model apparently believed it had already understood the image well enough and instead relied solely on text search or prior knowledge, thereby missing critical fine-grained visual cues. In the remaining case, the model did perform image search and retrieved relevant sources, but made a reasoning error in the final step.
These observations indicate that the drop in Easy-split accuracy is not primarily driven by over-retrieval or exposure to distractors, but by underuse of image search in the more complex full-rollout setting: the model sometimes “skips” a necessary image-search step when it incorrectly assumes that additional retrieval is unnecessary. We therefore attribute the Easy-split gap mainly to:
@@
Tool-use decision errors. The dominant failure mode for o3 is a policy-level issue—deciding whether to invoke image search—rather than over-retrieval or inherent drawbacks of the full-rollout setting.
Overall, rollout helped where progressive narrowing was necessary (academic, tech, media queries), but on easy items performance was limited by miscalibration in tool choice: the model occasionally declined to use image search even when it was crucial for success.
Adding SoM (Full Rollout + SoM) gave consistent, targeted gains.
The improvements were o3 37.6 (+1.6), Gemini 27.7 (+3.9), Qwen 7.1 (+1.0) over Full Rollout. Set-of-Mark anchored reasoning on fine-grained visual cues (micro-text, logos, key entities) before expanding the search space.
Decoupling internal knowledge.
To disentangle the effect of models’ internal knowledge from their use of external tools, we further evaluate on MMSearch, a subset of 239 tasks that are unsolvable without search by any model we tested (Table 4). The relationship between performance on the full and lite benchmarks is visualized in Figure 4.
On MMSearch, the best-performing setting is full rollout with only text search (abbreviated as Text Search in the table) with the o3 model, achieving 31.4% accuracy. This suggests that, once questions solvable purely from internal memory are removed, current MLLMs still struggle to exploit fine-grained image search effectively. In many categories, we observe a drop in accuracy when moving from only text search to full rollout / full rollout + SoM. However, there remain categories where image search is clearly beneficial: for o3 in Sports, accuracy increases by 6.6 (full rollout) and 10.0 points (full rollout + SoM) over text-only rollout. Overall, most methods lie close to the diagonal in Figure 4, indicating that the qualitative trends observed on MMSearch+ are not driven primarily by parametric memorization, but by genuine tool-use and retrieval behavior.
@@ 5.1 Observations
In Figure 5, all models made more text-search calls than image-search calls. We analyze the contribution of the “zoom-in then image-search” tool-call pattern to the final accuracy. For Gemini-2.5-Pro, out of 39 newly answered correctly questions from full rollout to full rollout + SoM, there are 5 such cases in which the model zoomed in on an image and then performed an image search (12.8%). For o3, out of 42 newly answered correctly questions, there is 1 such case (2.4%). This indicates that Gemini utilized fine-grained image search more frequently, while o3 often zoomed in simply to view a region more clearly.
To empirically support this observation, we examine the Markov transition probabilities between tool types (Figure 12). In particular, Set-of-Mark allows cropping and sub-image search, so we measure how often a zoom action is immediately followed by an image search. The transition probability varies substantially across models: Gemini uses image search after zooming in 25.37% of cases (34/134), Qwen in 10.56% (15/142), and o3 in only 2.87% (18/627). These statistics are consistent with the accuracy gains above—Gemini tends to use zoomed-in regions as retrieval queries, whereas o3 seldom does, often preferring further zooms or text search instead.
Image search and text search benefit different task categories.
The accuracies of o3 with text search (up to 10 rounds) and with full rollout are similar (37.0 vs. 36.0). However, their category-level accuracy distributions differ substantially. For example, image search yields better performance in Film/TV, whereas additional text search rounds provide greater benefits in Geography.
Scene-text is predominantly processed as textual evidence rather than visual evidence.
Across tasks containing logos, signage, or other forms of scene text, we observed that MLLMs overwhelmingly chose to convert visual text into explicit textual queries rather than invoking image search on the corresponding subregions. By detecting scene text with DeepSeek-OCR and manually inspecting sampled trajectories, we found that both o3 and Gemini-2.5-Pro typically relied on their built-in OCR capabilities to extract strings and issue text-search calls. For o3, 8 out of 10 sampled trajectories used only text search, while Gemini behaved similarly in 7 of 10 cases. Direct image-search calls on SoM-highlighted text regions were rare and generally appeared only when the model seemed uncertain about the exact transcription or wished to leverage additional visual cues such as layout or font. These findings suggest that, within MMSearch-Plus, scene-text reasoning is mediated primarily through textual grounding rather than visual retrieval, highlighting that multimodal agents implicitly choose between two tool pathways, (a) OCR + text search versus (b) image search, based on confidence and perceived utility.
@@ 6 Related Work
Classical vision–language benchmarks probe compositional and commonsense reasoning in images and text (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021). Among these, GQA (Hudson and Manning, 2019) emphasizes structured compositional reasoning using real-world scenes, while OK-VQA (Marino et al., 2019) requires external knowledge beyond the visual signal. More recently, InfoSeek (Chen, 2023) frames VQA as an information-seeking task, but relies on offline retrieval corpora whose contents frequently overlap with pretraining data for modern MLLMs, weakening its ability to measure genuine retrieval behavior. LiveVQA (Fu et al., 2025) benchmarks live visual knowledge by deriving VQA instances from recent web content beyond models’ cutoff dates, and shows that multimodal search augmentation substantially boosts performance.
Beyond static images, recent benchmark suites extend to long videos and interleaved modalities (Wu et al., 2024; Li et al., 2024; Hu et al., 2025; Jeong et al., 2025). Concurrent work on web agents studies planning, browsing, and evidence gathering in open worlds (Krishna et al., 2024; Wei et al., 2025; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025).
Multimodal web search benchmarks begin to couple vision with browsing but often emphasize entity matching or template-based grounding over fine-grained visual reasoning (Jiang et al., 2024; Wu et al., 2025). Closest to our setting, MM-BrowseComp targets image/video evidence encountered during search and reveals substantial headroom even for frontier models (Li et al., 2025b).
A complementary line promotes “thinking with images” via improved grounding or RL-style training to elicit image-first reasoning (Wu and Xie, 2024; Zheng et al., 2024; Fan et al., 2025; Zheng et al., 2025b; Su et al., 2025a; Xia et al., 2025).
We differ from prior benchmarks by targeting sparse, spatial–temporal visual cues that must be transformed into targeted queries and multi-step browsing plans. Crucially, tasks in MMSearch-Plus are designed to be extremely unlikely to be answerable from internal model knowledge alone, requiring models to (i) decide what visual regions matter, (ii) choose appropriate external tools (image vs. text search), and (iii) gather and verify evidence online. This provides a focused stress test for agentic multimodal browsing systems.
@@ 7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) achieved less than 40% accuracy, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieved 13.5%. These results indicate that current systems struggle to (i) decide when and how to crop and reuse visual evidence, and (ii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Ethics Statement
Principles.
We follow the Code of Ethics: uphold scientific rigor, avoid harm, ensure transparency, foster fairness, respect intellectual labor, and protect privacy and confidentiality.
Data sources, consent, and licensing.
We curate only publicly accessible material from (i) consumer video platforms (keyword queries) and (ii) open-access repositories (e.g., arXiv). We comply with platform Terms of Service and licenses and preserve attributions (e.g., video URL, paper URL). Items are treated as “events” and used solely for research.
Privacy, minimization, and redaction.
We mask sensitive content (e.g., faces, license plates); we exclude minors and content revealing private spaces when feasible. Residual risks remain; we provide an appeals/takedown channel.
Fairness.
To counter skews in public platforms, we conducted manual sampling across predefined categories and regions, report measurable distributions, and forbid uses that target protected classes or amplify inequities.
Limitations.
While diverse, our data inherits biases from public web sources and the domains we sampled; some tasks still under-represent non-English or low-visual-density pages. Our evaluation focuses on general-purpose agents with standard tool access; specialized systems with richer UI control or domain knowledge may behave differently. Finally, we primarily target image–text browsing; videos and dynamic interfaces remain underexplored.
Reproducibility Statement
We have taken several steps to facilitate independent verification of our results. Implementation details, hyperparameters, and evaluation protocols are documented in Section 3. We uploaded a randomly selected subset (seed = 42, 30 task examples) as supplementary material due to the 100MB space constraint, and released an anonymized archive on how to use the dataset in https://anonymous.4open.science/r/MMSearch-Plus-D54E.